# Computing Large 2D/3D Convex Regions of Obstacle-Free Space for Robotic Planning

### v2

@@ III-A Problem Formulation
We require no collision between the seed $\mathcal{Q}$ and the obstacle region $\mathcal{O}$, thus implying $\mathcal{Q}\cap\mathcal{O}=\varnothing$.
@@
$\max_{\mathcal{P}}~{}{\operatorname*{vol}({\mathcal{P}})},~{}~{}\mathrm{s.t.}~{}\mathcal{Q}\subseteq\mathcal{P},~{}\mathcal{O}\subseteq\operatorname*{coint}({\mathcal{P}}),$ (3)
where $\operatorname*{vol}({\mathcal{P}})$ denotes the volume of the convex polytope $\mathcal{P}$ and $\operatorname*{coint}({\cdot})$ denotes the complement of the interior of a set. Note that the solution set of (3) will never be empty since the seed $\mathcal{Q}$ itself is already a feasible solution.
@@ III-B Algorithm Overview
Even if we simplify the optimization (3) by disregarding the constraint of including the seed, specifically $\mathcal{Q}\subseteq\mathcal{P}$, it will be transformed into a generalization of the challenging potato peeling problem, which has an $O(N^{7})$ approach [39] in 2-D case and is NP-hard [40] in 3-D case. Not to mention that bringing in this constraint complicates the optimization (3) even more than potato peeling. To achieve a balance between efficiency and quality, we only seek a high-quality feasible solution to (3) efficiently instead of the global optimal solution. Additionally, since the computation of the objective function vol$(\mathcal{P})$ of (3) is at least as hard as an NP-complete problem [41], we optimize a reasonable lower bound of vol$(\mathcal{P})$, which is straightforward to evaluate, to maximize the original objective function. As adopted in [1], we also choose the the volume of the MVIE of $\mathcal{P}$ as this lower bound.
@@ 7             $\mathcal{E}\leftarrow\mathcal{E}_{k-1},~{}\bar{\mathcal{P}}\leftarrow\mathbb{R}^{n}$
11                   $a_{i}\leftarrow\operatorname*{arg\,max}_{a\in\mathbb{R}^{n}}{a^{\mathrm{T}}a},$ ${}~{}~{}~{}~{}~{}~{}\mathrm{s.t.}~{}\bar{\mathcal{Q}}\subseteq\mathcal{H}({a}),~{}\bar{\mathcal{O}}_{i}\subseteq\operatorname*{coint}({\mathcal{H}({a})})$
@@ 16                   $\bar{\mathcal{P}}\leftarrow\bar{\mathcal{P}}\cap\mathcal{H}({a_{j}})$
17                   \$\mathcal{I}\leftarrow\mathcal{I}\backslash\left\{{i\in\mathcal{I}~{}\big{|}~{}\bar{\mathcal{O}}_{i}\subseteq\operatorname*{coint}({\mathcal{H}({a_{j}})})}\right\}$
@@ III-B1 RsI
where $a_{i}$ is the only contact point of the boundary of the halfspace with the inflated ball $\mathcal{B}_{i}$. Then the halfspace $\mathcal{H}({a_{i}})$ can be defined as
@@
The calculation of the restrictive halfspace can be written as an optimization as shown in Line 1, for which we provide an efficient solution method in Sec. IV. In the left diagram of Fig. 2(a), we provide a specific example of the restrictive halfspace computation in the transformed space.
In the second inner loop (Line 1-1), we generate a new convex polytope based on the obtained halfspaces. We iteratively finds the closest halfspace $\mathcal{H}({a_{i}})$, adds it into $\bar{\mathcal{P}}$, and then remove the halfspaces corresponding to the obstacles outside of $\mathcal{H}({a_{i}})$. Until all halfspaces are processed, an obstacle-free polytope in $\mathcal{H}$-representation [38] can be formed in the transformed space. Then a new polytope $\mathcal{P}_{k}$ is obtained by recovering $\bar{\mathcal{P}}$ to the original space (Line 1). Thus, the output of FIRI is an obstacle-free convex polytope in $\mathcal{H}$-representation, i.e., intersection of $m$ halfspaces
@@ III-B2 MVIE
which is employed in Line 1. To solve MVIE, we propose efficient methods in Sec. V and Sec. VI so that the computational overhead of MVIE will no longer be a stumbling block that prevents this monotonically inflating MVIE approach from being applied to real-time scenarios [1].
@@ III-B3 Manageability and Convergence of FIRI
$\mathcal{Q}\subseteq\mathcal{P}_{k},~{}\mathcal{O}\subseteq\operatorname*{coint}({\mathcal{P}_{k}}),$ (12)
@@
In conclusion, RsI brings manageability to FIRI, and the monotonically iterative updates allows FIRI to continuously optimize the lower bound of the polytope’s volume to obtain a high quality output. It is worth emphasizing that the efficiency of FIRI relies strongly on the performance of solving two optimization programming in Line 1 and 1. In subsequent sections, we propose novel subalgorithms that are practically efficient and reliable, exploiting the geometrical structure of the involved programming, which greatly benefits the computational efficiency of FIRI.
@@ IV-A Reformulation of Restrictive Halfspace Computation
Algorithm 1 keeps $a_{i}^{\mathrm{T}}a_{i}>0$, which ensures that the origin always lies within the interior of the the halfspace $\mathcal{H}({a_{i}})$. Thus, although such a maximization of the inflated ball $\mathcal{B}_{i}$ is nonconvex, we can obtain its equivalent minimum-norm formulation through its polar duality. Specifically, we directly optimize the polar dual vector of $a_{i}$ by substituting $a_{i}=b/(b^{\mathrm{T}}b)$ and obtain an equivalent $L_{2}$-norm minimization
@@
which has a low-dimension but multi-constraint nature.
@@ IV-B2 Recursive Problem Construction
An essential process of Algorithm 2 is the recursive problem construction (Line 2-2), which transforms (17) that has an equality constraint into a $(n-1)$ dimensional subproblem with the same form as the original problem (16). The efficiency and numerical stability of this process have a great impact on Algorithm 2. Since $L_{2}$-norm is invariant under the orthogonal transformation, we establish a new $(n-1)$ dimensional Cartesian coordinate system on the constraint plane (17c) to implement the construction of the reduced dimensional recursive problem. We provide a specific example of this process in Fig. 3(c) and 3(d). As illustrated in Fig. 3(c), we take the point of $L_{2}$-norm minimization on the constraint plane as the origin of the coordinate system,
@@
where $e_{j}\in\mathbb{R}^{n}$ is an unit vector whose $j$-th element is 1. We provide the details of the choice of $j$ and the computation of $H$ at the end of Sec. IV-B2. The requirement (20) indicates that the set of all column vectors of the orthogonal $H^{\mathrm{T}}$ except the $j$-th one form an orthonormal basis for the orthogonal complement of $v$, as shown in Fig. 3(c). Since $v$ is normal to the constraint plane, we adopt this orthonormal basis to define the new coordinate system of the $(n-1)$ dimensional subspace on the constraint plane, and introduce the corresponding coordinates $y^{\prime}\in\mathbb{R}^{n-1}$. We denote $M\in\mathbb{R}^{n\times(n-1)}$ as the matrix obtained by removing the $j$-th column of $H^{\mathrm{T}}$. Then $y$ in the origin coordinate system corresponding to $y^{\prime}$ can be computed as
@@
According to the decomposition (19) and (21), minimizing the $L_{2}$-norm of $y$ on the constraint plane is equivalent to minimizing
@@ IV-B3 Complexity Analysis
To conclude this section, we give the complexity analysis of the randomized Algorithm 2. There is a probability to trigger a recursive call of the $(n-1)$ dimensional subproblem in each iteration. The idea of this randomized algorithm is that when the dimension $n$ of the problem is small but the constraint scale $d$ is huge, the probability of a newly added constraint becoming an active constraint is tiny, thus the probability of recursive calls is tiny as well. The expected complexity of this algorithm is $O(n!d)$ [13], so that in the common dimensions $n\in\{2,3\}$, the expected complexity only increases linearly with the constraint scale. In addition, since the algorithm preprocess the order of constraints by a random permutation, its actual performance does not depend on the input order and almost always achieves expected linear complexity. With the help of this randomized algorithm, the complexity of handling all obstacles in the RsI of Algorithm 1 grows linearly only with the total number of vertices of the obstacles and the seed.
@@ V Solving SOCP-Reformulation of MVIE via Affine Scaling Algorithm
In this section, we propose a practically efficient algorithm for MVIE in a low dimension but with a large number of constraint, which is required in FIRI (Line 1 of Algorithm 1). By carefully handling the orthogonal constraints on the elliptical matrix and the objective function involving the matrix determinant, we reformulate the MVIE problem into a Second-Order Conic Programming (SOCP) form. Then we employ Affine Scaling [45] algorithm for solving this SOCP, achieving efficient solution of MVIE.
@@ V-A SOCP-Reformulation of MVIE
where $m$ denotes the number of halfspaces that compose the convex polytope $\mathcal{P}$ defined in (10). Additionally, $\mathcal{K}_{1/n}$ can also be represented by SOC using additional $O(n)$ variables and cones of $\mathcal{K}_{3}$ [20]. Eventually, we transform (34) into an optimization with only SOC form constraints and a simple linear objective function.
@@ V-B Affine Scaling for Solving SOCP with Massive Constraints
where $x\in\mathbb{R}^{\bar{n}}$, $c_{\mathcal{K}}\in\mathbb{R}^{\bar{n}}$, ${c}_{i}\in\mathbb{R}^{\bar{n}}$, ${d}_{i}\in\mathbb{R}$ and ${A}_{i}$ are all constant. The new decision variable $x$ consists of all lower triangular elements of $L_{\mathcal{E}}$, $b_{\mathcal{E}}$, $t$ of (34), and the elements added in order to deal with constraint (34b) in form of hypograph of geometric mean. We have $\bar{n}=O(n^{2})$ because an $n$ dimensional ellipsoid already has $n(n+3)/2$ variables. For convenience, we set $t$ to be the $\bar{n}$-th element of $x$, then we have
@@
$\bar{m}=O(m+n)$ indicates the amount of the SOC constraints. For the constraints in (35b) corresponding to the origin constraints of (34b) which are represented by additional cones of $\mathcal{K}_{3}$, ${A}_{i}\in\mathbb{R}^{\bar{n}\times 2}$ and $n_{i}=3$. For the constraints in (35b) corresponding to the origin constraints of (34c), ${A}_{i}\in\mathbb{R}^{\bar{n}\times n}$ and $n_{i}=n+1$.
@@ VI Solving 2-D MVIE Analytically via A Linear-time Complexity Algorithm
Inspired by the linear-time algorithm of the dual problem Minimum Volume Enclosing Ellipsoid (MVEE) [24], in this section, we focus on the construction of an algorithm that achieves linear-time complexity for 2-D MVIE, leveraging its special LP-type problem [25] structure. We begin by providing the necessary background knowledge related to LP-type problems and analyzing the limitations of existing general solutions for applying to MVIE in Sec. VI-A. Then we address these limitations and propose our algorithm with linear time complexity in Sec. VI-B. Finally we provide the analytic solution of the subproblems required in the proposed algorithm in Sec. VI-C.
@@ VI-B1 Value Function
We design an additional evaluation of the proximity to closure for the unclosed case and use the MVIE volume for the closed case to construct a new value function that satisfies the properties of the LP-type problem. We define the value domain $\mathcal{W}$ in lexicographical ordering set $(\mathcal{W},<)$. We denote $<$ as the lexicographical ordering on $\mathbb{R}^{2}$, that is, for $\forall x=(x_{1},x_{2})^{\mathrm{T}},y=(y_{1},y_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$, $x<y$, if $x_{1}<y_{1}$, or $x_{1}=y_{1}$ and $x_{2}<y_{2}$. This ordering is extended to $\mathbb{R}^{2}\cup\{-\infty\}$ by the definition that $\infty$- ∞ is the unique minimum value. Considering the frequent failure to form closed polygon for the subsets of constraints in MVIE, we define the value function as
@@ VI-B2 Basis Computation
where $\mathcal{P}$ is the input polygon and also denotes the set of the finite halfspaces whose intersection forms the input polygon in this section. ${\bar{\mathcal{P}}^{N}}$ is its subset containing $N$ elements. Additionally, ¯$\bar{\mathcal{E}}(\bar{\mathcal{P}}^{N})$ is the subproblem defined by
@@
Since at least $3$ halfspaces are required to form a closed space in 2-D and an ellipse has only $5$ degrees of freedom, we require $3\leq N\leq 5$. When ${\bar{\mathcal{P}}^{N}}$ is closed and has no redundant element, (48) has an analytical solution, and its result $\bar{\mathcal{E}}$ is the Maximal Ellipse tangent to N edges of the N-gon formed by ${\bar{\mathcal{P}}^{N}}$, denoted as MENN, which is detailed in Sec. VI-C. Note that $\mathcal{E}^{*}$ and $\bar{\mathcal{E}}$ are used to denote MVIE and MENN, respectively. For the two cases where the subproblems (48) are infeasible, we handle them with special treatment: i) If ${\bar{\mathcal{P}}^{N}}$ is not closed, we define its $\bar{\mathcal{E}}$ as $\infty$. ii) If there are redundant elements in a closed ${\bar{\mathcal{P}}^{N}}$, that is, these elements are not any side of the formed closed polygon. At this point, it is impossible for these elements to satisfy (48c), then we also define its $\bar{\mathcal{E}}$ as $\infty$. Based on the subproblem (48), MVIE (47) can be solved through enumerating subproblems (48) which has analytical solution.
@@
where $\bar{\mathcal{P}}^{3}_{j}$ denotes the subset, which contains $3$ elements, of $\bar{\mathcal{P}}^{4}_{1}$. This equation (49) corresponds to the process in the large orange box in Fig. 4: among the MVIE for each subset $\bar{\mathcal{P}}_{j}^{3}$, and the MENN of $\bar{\mathcal{P}}^{4}_{1}$, we select the largest one that satisfies the constraint (49b) as the MVIE of $\bar{\mathcal{P}}_{1}^{4}$. Based on the result we can build upwards MVIE of $\mathcal{P}^{5}$ containing $5$ elements in a similar way, as illustrated in the blue box of Fig. 4. Eventually, for the input set $\mathcal{P}$ containing more than $5$ elements, its MVIE can be obtained by calculating the MVIE of all its subsets $\bar{\mathcal{P}}^{5}$ and selecting the largest ellipse contained within $\mathcal{P}$. Additionally, it should be noted that the MENN of a closed $\bar{\mathcal{P}}^{3}_{j}$ is its MVIE  [49]. To summarize, we can now compute the MVIE of the input set $\mathcal{P}$ based entirely on its subsets’ MENN ($3\leq N\leq 5$) which can be solved analytically in an efficient enumeration order.

### v3

@@ III-A Problem Formulation
which additionally enables FIRI to directly process polytope-type obstacles, as opposed to point-only methods [2, 9] that require discretization to handle such obstacles. We require no collision between the seed $\mathcal{Q}$ and the obstacle region $\mathcal{O}$, thus implying $\mathcal{Q}\cap\mathcal{O}=\varnothing$.
@@
$\max_{\mathcal{P}}~{}{\operatorname*{vol}({\mathcal{P}})},~{}~{}\mathrm{s.t.}~{}\mathcal{Q}\subseteq\mathcal{P},~{}{\color[rgb]{0,0,0}\mathcal{O}\cap\text{int}({\mathcal{P}})=\varnothing},$ (3)
where $\operatorname*{vol}({\mathcal{P}})$ denotes the volume of the convex polytope $\mathcal{P}$ and $\text{int}({\cdot})$ denotes the interior of a set. Note that the solution set of (3) will never be empty since the seed $\mathcal{Q}$ itself is already a feasible solution.
@@ III-B Algorithm Overview
Even if we simplify the optimization (3) by disregarding the constraint of including the seed, specifically $\mathcal{Q}\subseteq\mathcal{P}$, it will be transformed into a challenging problem, requiring $O(N^{7})$ [42] in 2-D and NP-hard [43] in 3-D. Not to mention that bringing in this constraint further complicates the optimization (3). For efficiency-quality balance, we only seek a high-quality feasible solution to (3) efficiently instead of the global optimal solution. Additionally, since the computation of the objective function vol$(\mathcal{P})$ of (3) is at least as hard as an NP-complete problem [44], we optimize a reasonable lower bound of vol$(\mathcal{P})$, which is straightforward to evaluate, to maximize the original objective function. As adopted in [1], we also choose the volume of the MVIE of $\mathcal{P}$ as this lower bound.
@@ 7             $\mathcal{E}\leftarrow\mathcal{E}_{k-1},~{}\bar{\mathcal{P}}\leftarrow\mathbb{R}^{n}$
11                   $a_{i}\leftarrow\operatorname*{arg\,max}_{a\in\mathbb{R}^{n}}{a^{\mathrm{T}}a},$ ${}~{}~{}~{}~{}~{}~{}\mathrm{s.t.}~{}\bar{\mathcal{Q}}\subseteq\mathcal{H}({a}),~{}{\color[rgb]{0,0,0}\bar{\mathcal{O}}_{i}\cap{\text{int}({\mathcal{H}({a})})}=\varnothing}$
@@ 16                   $\bar{\mathcal{P}}\leftarrow\bar{\mathcal{P}}\cap\mathcal{H}({a_{j}})$
17                   \$\mathcal{I}\leftarrow\mathcal{I}\backslash\left\{{i\in\mathcal{I}~{}\big{|}~{}{\color[rgb]{0,0,0}\bar{\mathcal{O}}_{i}\cap{\text{int}({\mathcal{H}({a_{j}})})}=\varnothing}}\right\}$
@@ III-B1 RsI
where $a_{i}$ is the only contact point of the halfspace boundary with the inflated ball $\mathcal{B}_{i}$. The halfspace $\mathcal{H}({a_{i}})$ is defined as
@@
The calculation of the restrictive halfspace can be written as an optimization as shown in Line 1, for which we provide an efficient solver in Sec. IV. In the left diagram of Fig. 2(b), we provide a specific example of the restrictive halfspace computation in the transformed space, comparing scenarios with and without manageability. Notably, in the scenario without manageability (employed by IRIS), the exclusive focus on maximizing ellipsoid inflation results in the computed halfspace being unable to guarantee seed point containment.
In the second inner loop (Line 1-1), we generate a new convex polytope based on the obtained halfspaces. We iteratively finds the closest halfspace $\mathcal{H}({a_{i}})$, adds it into $\bar{\mathcal{P}}$, and then remove the halfspaces corresponding to the obstacles outside of $\mathcal{H}({a_{i}})$. This widely adopted greedy strategy [1, 2, 31] is effective in reducing the number of halfspaces that constitute the generated convex polytope, aiming to minimize them as much as possible, compared to the large number of obstacles. Until all halfspaces are processed, an obstacle-free polytope in $\mathcal{H}$-representation [41] can be formed in the transformed space. Then a new polytope $\mathcal{P}_{k}$ is obtained by recovering $\bar{\mathcal{P}}$ to the original space (Line 1). Thus, the output of FIRI is an obstacle-free convex polytope in $\mathcal{H}$-representation, i.e., intersection of $m$ halfspaces
@@ III-B2 MVIE
which is employed in Line 1. Notably, MVIE takes exclusively one convex polytope as input, independent of any obstacles used to generate the input polytope. To solve MVIE, we propose efficient methods in Sec. V and Sec. VI so that the computational overhead of MVIE will no longer be a stumbling block that prevents this monotonically inflating MVIE approach from being applied to real-time scenarios [1].
@@ III-B3 Manageability and Convergence of FIRI
$\mathcal{Q}\subseteq\mathcal{P}_{k},~{}{\color[rgb]{0,0,0}\mathcal{O}\cap\text{int}({\mathcal{P}_{k}})=\varnothing}$ (12)
@@
In conclusion, RsI brings manageability to FIRI, while monotonic iterative updates optimize the polytope volume’s lower bound for a high-quality output. Notably, the efficiency of FIRI relies strongly on the performance of solving two optimizations in Line 1 and 1. In subsequent sections, we propose efficient and reliable subalgorithms exploiting the geometric structure of these optimizations, greatly benefiting the computational efficiency of FIRI.
@@ IV-A Reformulation of Restrictive Halfspace Computation
Algorithm 1 keeps $a_{i}^{\mathrm{T}}a_{i}>0$, which ensures that the origin always lies within the interior of the halfspace $\mathcal{H}({a_{i}})$. Thus, although such a maximization of the inflated ball $\mathcal{B}_{i}$ is nonconvex, we can obtain its equivalent minimum-norm formulation through its polar duality. Specifically, we reformulate the problem with a new variable $b$ by substituting $a_{i}=b/(b^{\mathrm{T}}b)$ and obtain an equivalent $L_{2}$-norm minimization
@@
which has a low-dimension but multi-constraint nature. In reformulation (15), both types of constraints share a unified inequality form. This formulation handles cases with and without the seed containment constraint, allowing subsequent solver SDMN adapt readily. In contrast, prior work such as IRIS [1] struggles to add seed containment while maintaining the original problem structure, making it challenging to achieve manageability via straightforward constraint addition.
@@ IV-B2 Recursive Problem Construction
An essential process of Algorithm 2 is the recursive problem construction (Line 2-2), which transforms (17) with an equality constraint into a $(n-1)$ dimensional subproblem with the same form as the original problem (16). This process’s efficiency and stability greatly impact Algorithm 2. Since $L_{2}$-norm is invariant under the orthogonal transformation, we establish a new $(n-1)$ dimensional Cartesian coordinate system on the constraint plane (17c) to implement the construction of the reduced dimensional recursive problem. We provide a specific example of this process in Fig. 3(c) and 3(d). As illustrated in Fig. 3(c), we take the point of $L_{2}$-norm minimization on the constraint plane as the origin of the coordinate system,
@@
where $e_{j}\in\mathbb{R}^{n}$ is an unit vector whose $j$-th element is 1. We provide the details of the choice of $j$ and the computation of $H$ at the end of Sec. IV-B2. The requirement (20) indicates that the set of all column vectors of orthogonal $H^{\mathrm{T}}$ except the $j$-th one form an orthonormal basis for the orthogonal complement of $v$, as shown in Fig. 3(c). Since $v$ is normal to the constraint plane, we adopt this orthonormal basis to define the new coordinate system of the $(n-1)$ dimensional subspace on the constraint plane, and introduce the corresponding coordinates $y^{\prime}\in\mathbb{R}^{n-1}$. We denote $M\in\mathbb{R}^{n\times(n-1)}$ as the matrix obtained by removing the $j$-th column of $H^{\mathrm{T}}$. Then $y$ in the origin coordinate system corresponding to $y^{\prime}$ can be computed as
@@
Based on the decomposition (19) and (21), minimizing the $L_{2}$-norm of $y$ on the constraint plane is equivalent to minimizing
@@ IV-B3 Complexity Analysis
To conclude this section, we give the complexity analysis of the randomized Algorithm 2. Each iteration may trigger recursive calls to $(n-1)$ dimensional subproblems. For small dimension $n$ with large constraint scale $d$, new constraints rarely become active, making recursive calls tiny. The expected complexity is $O(n!d)$ [13], so that in the common dimensions $n\in\{2,3\}$, the expected complexity increases only linearly with the constraint scale [^1] For specific derivation regarding the complexity, readers can refer to Theorem 1 of Seidel’s work [13].. In addition, random constraint permutation preprocessing ensures performance independent of input order, nearly always achieving expected linear complexity. With the help of this randomized algorithm, the complexity of handling all obstacles in the RsI of Algorithm 1 grows linearly only with the total number of vertices of the obstacles and the seed.
@@ V Solving SOCP-Reformulation of MVIE via Affine Scaling Algorithm
In this section, we propose an efficient algorithm for MVIE with a low dimension but massive constraints, required in FIRI (Line 1 of Algorithm 1). By carefully handling the orthogonal constraints on the elliptical matrix and the objective function involving the matrix determinant, we reformulate MVIE into a Second-Order Conic Programming (SOCP) form. Using Affine Scaling [48], we solve this SOCP efficiently for MVIE.
@@ V-A SOCP-Reformulation of MVIE
where $m$ denotes the number of halfspaces composing $\mathcal{P}$ defined in (10). Additionally, $\mathcal{K}_{1/n}$ can also be represented by SOC using additional $O(n)$ variables and cones of $\mathcal{K}_{3}$ [20]. Eventually, we transform (34) into an optimization with only SOC form constraints and a simple linear objective function.
@@ V-B Affine Scaling for Solving SOCP with Massive Constraints
where $x\in\mathbb{R}^{\bar{n}}$, $c_{\mathcal{K}}\in\mathbb{R}^{\bar{n}}$, ${c}_{i}\in\mathbb{R}^{\bar{n}}$, ${d}_{i}\in\mathbb{R}$ and ${A}_{i}$ are all constant. The new decision variable $x$ consists of all lower triangular elements of $L_{\mathcal{E}}$, $b_{\mathcal{E}}$, $t$ of (34), and the elements added to deal with constraint (34b) in form of hypograph of geometric mean. We have $\bar{n}=O(n^{2})$ because an $n$ dimensional ellipsoid already has $n(n+3)/2$ variables. For convenience, we set $t$ as the $\bar{n}$-th element of $x$, then we have
@@
$\bar{m}=O(m+n)$ indicates the amount of the SOC constraints. For the constraints in (35b) corresponding to the constraints of (34b) which are represented by additional cones of $\mathcal{K}_{3}$, ${A}_{i}\in\mathbb{R}^{\bar{n}\times 2}$ and $n_{i}=3$. For the constraints in (35b) corresponding to the origin constraints of (34c), ${A}_{i}\in\mathbb{R}^{\bar{n}\times n}$ and $n_{i}=n+1$.
@@ VI Solving 2-D MVIE Analytically via A Linear-time Complexity Algorithm
Inspired by the linear-time algorithm of the dual problem Minimum Volume Enclosing Ellipsoid (MVEE) [24], in this section, we focus on the construction of a linear-time algorithm for 2-D MVIE, leveraging its special LP-type problem [25] structure. We begin by providing the necessary background knowledge related to LP-type problems and analyzing the limitations of existing general solutions for applying to MVIE in Sec. VI-A. Then we address these limitations and propose our algorithm with linear time complexity in Sec. VI-B. Finally we provide the analytic solution of the subproblems required in the proposed algorithm in Sec. VI-C.
@@ VI-B1 Value Function
We design an additional evaluation of the proximity to closure for unclosed case and use the MVIE volume for closed case to construct a new value function that satisfies the properties of the LP-type problem. We define the value domain $\mathcal{W}$ in lexicographical ordering set $(\mathcal{W},<)$. We denote $<$ as the lexicographical ordering on $\mathbb{R}^{2}$, that is, for $\forall x=(x_{1},x_{2})^{\mathrm{T}},y=(y_{1},y_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$, $x<y$, if $x_{1}<y_{1}$, or $x_{1}=y_{1}$ and $x_{2}<y_{2}$. This ordering is extended to $\mathbb{R}^{2}\cup\{-\infty\}$ by the definition that $\infty$- ∞ is the unique minimum value. Considering the frequent failure to form closed polygon for the subsets of constraints, we define the value function as
@@ VI-B2 Basis Computation
where $\mathcal{P}$ is the input polygon while simultaneously denoting the finite set of the halfspaces whose intersection constitutes the input polygon in this section. Its subset containing $N$ elements is denoted as ${\bar{\mathcal{P}}^{N}}$. To prevent potential misunderstanding, it should be emphasized that the subset ${\bar{\mathcal{P}}^{N}}$ is derived from $\mathcal{P}$, the sole input of MVIE, and is independent of the obstacles used to generate $\mathcal{P}$ in RsI. Additionally, ¯$\bar{\mathcal{E}}(\bar{\mathcal{P}}^{N})$ is the subproblem defined by
@@
Since at least $3$ halfspaces are required to form a closed space in 2-D and an ellipse has only $5$ degrees of freedom, we require $3\leq N\leq 5$. When ${\bar{\mathcal{P}}^{N}}$ is closed and has no redundant element, (48) has an analytical solution, and its result $\bar{\mathcal{E}}$ is the Maximal Ellipse tangent to N edges of the N-gon formed by ${\bar{\mathcal{P}}^{N}}$, denoted as MENN, detailed in Sec. VI-C. Notably, $\mathcal{E}^{*}$ and $\bar{\mathcal{E}}$ are used to denote MVIE and MENN, respectively. For the two cases where subproblems (48) are infeasible, we handle them with special treatment: i) If ${\bar{\mathcal{P}}^{N}}$ is unclosed, we define its $\bar{\mathcal{E}}$ as $\infty$. ii) If there are redundant elements in a closed ${\bar{\mathcal{P}}^{N}}$, that is, these elements are not any side of the formed closed polygon. At this point, it is impossible for these elements to satisfy (48c), then we also define its $\bar{\mathcal{E}}$ as $\infty$. Based on the subproblem (48), MVIE (47) can be solved through enumerating subproblems (48) which has analytical solution.
@@
where $\bar{\mathcal{P}}^{3}_{j}$ denotes a subset, which contains $3$ elements, of $\bar{\mathcal{P}}^{4}_{1}$. Notably, the MVIE of a closed $\bar{\mathcal{P}}^{3}_{j}$ is exactly its MENN [52]. This equation (49) corresponds to the process in the large orange box in Fig. 4: among the MVIE for each subset $\bar{\mathcal{P}}_{j}^{3}$, and the MENN of $\bar{\mathcal{P}}^{4}_{1}$, we select the largest one that satisfies the constraint (49b) as the MVIE of $\bar{\mathcal{P}}_{1}^{4}$. Based on the result we can build upwards MVIE of $\mathcal{P}^{5}$ containing $5$ elements in a similar way, as illustrated in the blue box of Fig. 4. It should be emphasized that for the actual input $\mathcal{P}$, which typically consists of more than $5$ halfspaces, its MVIE is also computed through a similar bottom-up enumeration: calculating the MVIE of all its subsets $\bar{\mathcal{P}}^{5}$ and selecting the largest ellipse contained within $\mathcal{P}$. To summarize, by connecting the above processes through the bottom-up strategy, now we can compute the MVIE of the input $\mathcal{P}$ based entirely on its subsets’ MENN ($3\leq N\leq 5$) which can be solved analytically in an efficient enumeration order.